# Speed Limit Reduction Enhances Urban Worker Safety: Evidence from a Decade of Traffic Incidents in Santiago, Chile

### v1

@@ 5 Results
Our analysis of work-related traffic crashes, collisions, or falls in Santiago, Chile, revealed several interesting patterns across temporal, spatial, and demographic dimensions. We will present these findings, starting with the representativeness of our data, followed by exploratory analyses, and concluding with the results of our regression model examining the impact of the speed limit law.
@@ 5.1 Data Representativeness
To assess how well our MUSEG dataset represents overall travel patterns in Santiago, we compared it to the EOD2012 survey data. The Spearman correlation coefficient between the origin-destination matrices of these two datasets was $r=0.69$ ($p<$ 0.001). This positive correlation suggests that about 69% of the variability in the ranks of one dataset is explained by the relationship with the ranks of the other. This indicates that the travel patterns observed in our MUSEG dataset largely align with the general commuting patterns in Santiago, providing confidence in the representativeness of our data.
@@ 5.2 Mode of Transportation and Event Severity
Our analysis of event severity by mode of transportation revealed notable differences (Figure 2.a). Motorcycles, pedestrians and cycles were associated with the most severe incidents (44.45, 42.08 and 23.24 prescribed medical leave days, respectively). This pattern underscores the vulnerability of these road users.
@@ 5.3 Temporal Patterns
Most incidents in our dataset occurred during the morning peak hours (50.47% of incidents, with a mean severity of 20.55 days of prescribed medical leave; see Figure 3.b). This concentration can be attributed to the convergence of many workers starting their day at similar times. Interestingly, while the morning peak saw the highest frequency of incidents, night incidents (3.82%) were associated with greater severity (32.41 days of prescribed medical leave).
Figure 4: Number of incidents per gender and age ranges.
@@ 5.4 Demographic Patterns
Our analysis revealed significant gender differences in incident patterns (Figure 4). Female workers accounted for 37.42% of incidents in the MUSEG data, closely mirroring their 37% representation in the overall insured worker population (Mutual de Seguridad, 2023). Similarly, the EOD2012 survey reports 36% of commuting trips by female workers.
Gender disparities were evident in mode choice (Figure 4): males were involved in 87.05% of motorcycle and 74.66% of cycle incidents, while females were involved in 56.57% of public transport incidents. These patterns echo gender differences observed in Santiago’s latest travel survey (SECTRA, 2015), contrasting the worker composition ascribed to MUSEG, where 63% is male and 37% is female.
Age-wise, the 25-35-year age group accounted for the most significant fraction of incidents across all modes of transportation (Figure 4). However, the age distribution varied by mode, highlighting the importance of considering age in our subsequent analyses.
Regarding incident severity, we found that male workers tended to experience more severe incidents, aligning with previous research (Bermúdez et al., 2018). Additionally, incident severity increased with age for both genders (Figure 5).
@@ 5.5 Spatial Patterns
When examining incident patterns by mode of transportation (Figure 7), we found interesting spatial associations. Cycle incidents clustered in zones with well-developed cycleways and significant poles of attraction for work-related trips. In contrast, motorized vehicles (e.g., cars) exhibited a stronger association with the city’s peripheral sectors. These areas serve either as trip generators or intermediate pass-through zones, potentially leading to higher vehicle speeds compared to other city areas. A key takeaway from these spatial patterns is the realization that infrastructure alone does not ensure safety, especially for cyclists.
@@ 5.6 Speed Limit Law Effects
Before proceeding with the regression, we compared the distribution of medical leave days between geolocated and non-geolocated incidents (Figure 8). The distributions were similar for incidents resulting in up to two years of medical leave. Consequently, we excluded the small fraction of incidents (0.001%) with longer leave periods from our analysis. We also excluded taxi trips (1.32% of the dataset) due to their distinct distribution pattern. In total, we discarded 1.24% of observations according to this criteria.
@@
We calculated the marginal effects at the mean for each variable to ease interpretation of the results (Figure 9). The marginal effect represents the change in the expected number of prescribed medical leave days for a one-unit change in the independent variable, holding all other variables constant at their mean values. For categorical variables, it represents the difference in expected days of leave when the variable is true versus false, or when compared to the reference category. A positive marginal effect implies more severe injuries, as it is associated with more prescribed days of medical leave.
@@
    Demographics: Female workers tended to have less severe injuries (7.57 fewer days of leave on average). This gender difference in injury severity has been observed in other studies (DeJoy, 1992; González-Sánchez et al., 2021; Havet et al., 2021). Age was also a significant factor, with each year associated with 0.57 additional days of leave, reflecting the increased vulnerability of older workers to severe injuries.
@@
    Speed Limit Law: Importantly, our analysis showed that the speed limit law was associated with a decrease of 4.26 days in prescribed medical leave for incidents in urban areas. This suggests that the law had a positive impact on reducing injury severity in work-related traffic incidents.
These results, visualized in Figure 9, provide valuable insights into the factors influencing work-related traffic incident severity in Santiago and the potential benefits of speed management policies in urban areas. The marginal effects allow us to quantify the relative impact of each factor on incident severity, providing a clear picture of which elements have the strongest influence on the length of medical leave following a work-related traffic incident.
@@ 6 Conclusions
Our analysis of work-related traffic incidents in Santiago, Chile, over a decade, revealed significant patterns in incident occurrence and severity across temporal, spatial, and demographic dimensions. These findings have important implications for urban planning, transportation policy, and workplace safety initiatives.
The study identified motorcycles, cycles, and pedestrian trips as modes associated with more severe injuries compared to motorized vehicles. This vulnerability of certain users highlights the need for targeted safety measures (Miner et al., 2024; Mangones et al., 2024). We observed gender differences in incident severity, with female workers experiencing less severe injuries on average. Older workers were more likely to suffer more severe injuries (DeJoy, 1992; González-Sánchez et al., 2021; Elrud et al., 2019). These demographic factors suggest potential benefits from age and gender-specific safety interventions.
@@
A key finding of our study is the impact of the 2018 urban speed limit reduction. This policy change was associated with a decrease in injury severity, with incidents requiring an average of 4.26 fewer days of medical leave. This result provides empirical support for the effectiveness of speed management policies in improving road safety (Elvik, 2012; Cleland et al., 2020; Milton et al., 2021). The observed increase in car usage in Santiago during the study period aligns with trends reported by Graells-Garrido et al. (2023a). Although more people are affected by the speed limit law due to increased car use, the safety benefits extend to all road users, including pedestrians and cyclists.
This research contributes to efforts improving urban mobility and safety. Guzman et al. (2020) developed Sustainable Mobility Plans for organizations in Bogotá, highlighting the need for tailored strategies to address congestion and enhance safety in urban areas. The results have practical implications for policymakers and urban planners. The positive impact of the speed limit reduction supports expanding such policies in urban areas. The higher injury severity for certain road users calls for infrastructure improvements and safety measures to protect vulnerable groups. Differences across demographic groups and periods suggest benefits from targeted safety campaigns and interventions.
Integrating road safety into broader urban design strategies aligns with ‘safe system’ approaches to urban mobility (Wegman et al., 2012). However, we found a higher association of incidents with improved cycling infrastructure, showing that infrastructure alone does not ensure safety. A comprehensive approach to urban mobility safety should combine infrastructure improvements with education, enforcement, and policy measures (Banister, 2008). Future urban planning should consider these complex interactions to create safer urban environments.
Our study has limitations. The data comes from a single social insurance provider, which may only partially represent some work-related incidents in Santiago. Our analysis does not include fatal incidents, which may have different patterns and risk factors. Future research could incorporate data from multiple sources and include fatal incidents for a more comprehensive analysis, as this work complements other studies that have focused on fatalities and number of deaths. Our analysis relied on geolocated incidents, although a relevant fraction of incidents were not geolocated. We analyzed a specific subset of geolocated incidents to ensure experimental validity; however, future work could deepen the analysis by exploring individual municipalities. In this case, all data could be used.

### v2

@@ 4 Results
Our analysis of work-related transportation incidents in Santiago, Chile, revealed several interesting patterns across temporal, spatial, and demographic dimensions. We present these findings, starting with the representativeness of our data, followed by exploratory analyses, and concluding with the results of our regression model examining the impact of the speed limit law.
@@ 4.1 Data Representativeness
To assess how well our MUSEG dataset represents overall travel patterns in Santiago, we compared it to the EOD2012 survey data.
According to EOD2012, there are 1.88 million workers living in Santiago, who, on an average working day, perform 13.95 million trips, of which 18.7% are to-work trips, 2% are for-work trips, and 46% are to-home trips (with a fraction of these presumably mirroring work trips). We analyze the correlation between their home locations and work trips with those from MUSEG data. The Spearman correlation coefficient between the municipal home locations was $r=0.85$ ($p<$ 0.001), and between the origin-destination matrices was $r=0.69$ ($p<$ 0.001). These positive correlations suggests that about 85% and 69% respectively of the variability in the ranks of one dataset is explained by the relationship with the ranks of the other. This indicates that geographical location of worker’s homes and where they commute to observed in our MUSEG dataset largely align with the general patterns in Santiago, providing confidence in the representativeness of our data.
@@ 4.2 Mode of Transportation and Event Severity
Our analysis of event severity by mode of transportation revealed notable differences (Figure 2.a). Motorcycles, pedestrians and cycles were associated with the most severe incidents (means of 44.45, 42.08 and 23.24 prescribed medical days of leave, respectively). This pattern underscores the vulnerability of these road users.
@@ 4.3 Temporal Patterns
Most incidents in our dataset occurred during the morning peak hours (50.47% of incidents, with a mean severity of 20.55 days of prescribed medical leave; see Figure 3.b). This concentration can be attributed to the convergence of many workers starting their day at similar times. Interestingly, while the morning peak saw the highest frequency of incidents, night incidents (3.82%) were associated with greater severity (32.41 days of prescribed medical leave). This high concentration of incidents during morning peak hours likely reflects the rigid nature of work start times compared to more flexible evening departures. The greater severity of night incidents can be attributed to reduced visibility, higher vehicle speeds due to lower traffic volumes, and increased worker fatigue during night shifts. Studies of severe traffic injuries in the EU have found distinct temporal patterns, with afternoon and winter months being common times for severe pedestrian and car occupant injuries (Aarts et al., 2016).
Figure 4: Fraction of incidents per gender and age ranges (column-normalized).
@@ 4.4 Demographic Patterns
Our analysis revealed significant gender differences in incident patterns (Figure 4). Women workers accounted for 37.42% of incidents in the MUSEG data, closely mirroring their 37% representation in the overall insured worker population (Mutual de Seguridad, 2023). Similarly, the EOD2012 survey reports 36% of commuting trips by women workers.
Gender disparities were evident in mode choice: men were involved in 87.05% of motorcycle and 74.66% of cycle incidents, while women were involved in 56.57% of public transport incidents. These patterns echo gender differences observed in Santiago’s latest travel survey (SECTRA, 2015), contrasting the worker composition ascribed to MUSEG, where 63% is men and 37% is women.
The analysis of incident distribution across age groups (Figure 4) reveals distinct modal patterns. For motorcycle incidents, the proportion decreases with age, from 0.28 in the 18–25 age group to 0.13 in the 55–65 group. Conversely, public transportation incidents increase with age, from 0.23 to 0.33 in the same age ranges. Similar age-related increases are observed for pedestrian (0.06 to 0.09) and motorized incidents (0.16 to 0.21, peaking at 0.28 in the 35–45 age group). These shifts suggest age-dependent changes in transportation mode choice, though a detailed analysis of this relationship is beyond the scope of this study.
Regarding incident severity, we found that men workers tended to experience more severe incidents (Figure 5). Similar gender differences in medical leave duration were found by Bermúdez et al. (2018). Additionally, incident severity increased with age for both genders. The detailed quantification of these demographic effects is presented in the regression analysis section. However, without exposure data (e.g., kilometers traveled by gender), we cannot make conclusions about relative risk levels across demographic groups. The patterns described reflect incident characteristics rather than differential risk rates.
@@ 4.5 Spatial Patterns
When examining incident patterns by mode of transportation (Figure 7), we found interesting spatial associations. Cycle incidents clustered in zones with well-developed cycleways and significant poles of attraction for work-related trips. In contrast, motorized vehicles (e.g., cars) exhibited a stronger association with the city’s peripheral sectors. These areas serve either as trip generators or intermediate pass-through zones, potentially leading to higher vehicle speeds compared to other city areas. Using log-odds ratios, we found cycle incidents were more strongly associated with zones that have well-developed cycleways and significant poles of attraction for work-related trips, though our analysis does not assess the relationship between infrastructure quality and safety. While this analysis shows relative concentrations of incidents, it cannot determine actual risk rates without exposure data (trips or kilometers traveled by mode in each zone).
@@ 4.6 Speed Limit Law Effects
Before proceeding with the regression, we compared the distribution of medical days of leave between geolocated and non-geolocated incidents (Figure 8). The distributions were similar for incidents resulting in up to two years of medical leave. Consequently, we excluded the small fraction of incidents (0.001%) with longer leave periods from our analysis. We also excluded taxi trips (1.32% of the dataset) due to their distinct distribution pattern. In total, we discarded 1.24% of observations according to this criteria.
@@
We calculated the marginal effects at the mean for each variable to ease interpretation of the results (Figure 9). The marginal effect represents the change in the expected number of prescribed medical days of leave for a one-unit change in the independent variable, holding all other variables constant at their mean values. For categorical variables, it represents the difference in expected days of leave when the variable is true versus false, or when compared to the reference category. A positive marginal effect implies more severe injuries, as it is associated with more prescribed days of medical leave.
@@
    Demographics: Women workers tended to have less severe injuries (7.57 fewer days of leave on average). This gender difference in injury severity has been observed in other studies (DeJoy, 1992; González-Sánchez et al., 2021; Havet et al., 2021). Age was also a significant factor, with each year associated with 0.57 additional days of leave, reflecting the increased vulnerability of older workers to severe injuries.
@@
    Speed Limit Law: Importantly, our analysis showed that the speed limit law was associated with a decrease of 4.26 days in prescribed medical leave for incidents in urban areas. This suggests that the law had a positive impact on reducing injury severity in work-related transportation incidents.
These results, visualized in Figure 9, provide valuable insights into the factors influencing injury severity in Santiago and the potential benefits of speed management policies in urban areas. The marginal effects allow us to quantify the relative impact of each factor on injury severity, providing a clear picture of which elements have the strongest influence on the length of medical leave following a work-related transportation incident.
@@ 5 Conclusions
Our analysis of work-related transportation incidents in Santiago, Chile, over a decade, revealed significant patterns in incident occurrence and severity across temporal, spatial, and demographic dimensions. These findings have important implications for urban planning, transportation policy, and workplace safety initiatives.
A key finding of our study is the impact of the 2018 urban speed limit reduction. This policy change was associated with a decrease in injury severity, with incidents requiring an average of 4.26 fewer days of medical leave. This result provides empirical support for the effectiveness of speed management policies in improving road safety (Elvik, 2012; Cleland et al., 2020; Milton et al., 2021). The observed increase in car usage in Santiago during the study period aligns with trends reported by Graells-Garrido et al. (2023a). This finding is particularly significant as it provides evidence of speed limit effectiveness in a Latin American context, where such evaluations are scarce. While our results show an overall reduction in injury severity after the speed limit change, we cannot fully determine if the risk per user decreased for each mode of transportation. Changes in mode share during our study period mean that apparent reductions in incidents for certain modes could mask increased per-user risk if there were fewer total users of that mode. This is particularly relevant for vulnerable road users like pedestrians and cyclists.
The study identified motorcycles, cycles, and pedestrian trips as modes associated with more severe injuries compared to motorized vehicles. This vulnerability of certain users highlights the need for targeted safety measures (Miner et al., 2024; Mangones et al., 2024). We observed gender differences in incident severity, with women workers experiencing less severe injuries on average. Older workers were more likely to suffer more severe injuries (DeJoy, 1992; González-Sánchez et al., 2021; Elrud et al., 2019). These demographic factors suggest potential benefits from age and gender-specific safety interventions.
@@
This research demonstrates the effectiveness of speed management policies in improving urban mobility and safety. Guzman et al. (2020) developed Sustainable Mobility Plans for organizations in Bogotá, highlighting the need for tailored strategies to address congestion and enhance safety in urban areas. The results have practical implications for policymakers and urban planners. The positive impact of the speed limit reduction supports expanding such policies in urban areas. The higher injury severity for certain road users calls for infrastructure improvements and safety measures to protect vulnerable groups. While we observed differences in incident patterns across demographic groups, additional research incorporating exposure data (e.g., from contemporary travel surveys) would be needed to understand differential risks and design targeted interventions.
Integrating road safety into broader urban design strategies aligns with ‘safe system’ approaches to urban mobility (Wegman et al., 2012). A comprehensive approach to urban mobility safety should combine infrastructure improvements with education, enforcement, and policy measures (Banister, 2008). Future urban planning should consider these complex interactions to create safer urban environments.
Our study has limitations. The data comes from a single insurance provider, which may only partially represent some work-related incidents in Santiago. Our dataset lacks exposure data (such as kilometers traveled by different demographic groups), which limits our analysis of differential risk patterns. Our analysis does not include fatal incidents, which may have different patterns and risk factors. Furthermore, our dataset did not capture information about other vehicles involved in multi-vehicle incidents, preventing analysis of how different vehicle types contribute to injury severity. Studies have shown that larger vehicles pose disproportionate risks to other road users (Aldred et al., 2021), suggesting this is a crucial aspect for understanding traffic safety. Future research could incorporate data from multiple sources, include fatal incidents, and examine vehicle interaction patterns for a more comprehensive analysis. Our analysis relied on geolocated incidents, although a relevant fraction of incidents were not geolocated. We analyzed a specific subset of geolocated incidents to ensure experimental validity; however, future work could deepen the analysis by exploring individual administrative regions. In this case, all data could be used. Additionally, our spatial analysis cannot account for exposure levels by mode, preventing assessment of actual risk rates in different areas.